OpenStudy (anonymous):

63X^15y^9 ---------- 7xy^11 radical over the whole problem some one help?

OpenStudy (watchmath):

\(\sqrt{\frac{63x^{15}y^9}{7xy^11}}=\sqrt{\frac{9x^{14}}{y^2}}=\frac{3x^7}{y}\)

OpenStudy (anonymous):

thank you but could you please explain how you got that because i do know how to factor expressions but i cant do this one

OpenStudy (watchmath):

You do 63/7 to get 9 Then x^(15)/x to get x^14 Then y^9/y^(11) to get 1/y^2

OpenStudy (anonymous):

im still not foloowing you how do you ger x^14 from x^15 and how do you get 1Y^2 from y^11

OpenStudy (watchmath):

remember that x^m/x^n = x^(m-n) So x^(15)/x = x^(15-1)=x^(14) The others are similar

OpenStudy (anonymous):

oh i didint know that but thanks i think i get it